in 4 years Rs. 6000 amounts to Rs.8000.in what time will Rs. 525 amounts to 700 at the same ratte

Answer by mananth(16946)

On Rs 6000 interest = 2000

SI = P N R /100
2000=6000*4*r/100
2000*100/4*6000 = r
r=200/24
r= 8.33 %
Interest = 700-525= 175
P=525
I=175
r=8.33%
175= 525*N*8.33/100
N= 175*100/525*8.33
N=4 years